OpenStudy (anonymous):

Astronomers can use the brihtness of two light sources,such as stars, to compare the distances from the light sources.The intensity , or brightness, of light I is inversely proportionnal to the square of the distance from the light source d. a) rite an qution that represents this sititution. b) if d is the independent variable and I is the dependent variable,graph the rqution from exercise a when k=16 c) if two people are viewing the same light source,and one person is three time the distance from the light source as the other person,compare the light intensities t

OpenStudy (tad1):

a. I = k/d^2 b. Substitute values for d, make a table of the values; then use the table to plot the graph c. I1 = k/d; I2 = K/3d. the intensity of the second viewer will be 1/3 that of the first viewer.

OpenStudy (anonymous):

tad1 come back help

OpenStudy (tad1):

How can I help?

OpenStudy (anonymous):

can u show me made table

OpenStudy (tad1):

choose a value for d, any value as x is the independent variable. So I choose to be 1 Then I = K/1. If k is 16 then I = 16. Choose another value for d, say 5. then I = k/5=16/5 etc. Make a table for each of the intensities, graph the points and draw a rough graph.

OpenStudy (anonymous):

\[c) I_{1}=\frac{k}{d}\]\[12=\frac{k}{3d}\]?

OpenStudy (tad1):

Choose a value for d, not for I

OpenStudy (anonymous):

\[b)I=\frac{k}{1},\]

OpenStudy (tad1):

Yes. Do you know k? Pick other points for d such as d = 4. Then I = k/4, etc. Find enough points so that you can sketch the curve.

OpenStudy (anonymous):

ty